OpenStudy (anonymous):

HELP I WILL FAN AND MEDAL Complete the function table. Then write a rule for the function. Input output -3 -2 -4 -1 0 1 -1

OpenStudy (anonymous):

i know the first out put is -9

OpenStudy (anonymous):

-5 -3 -2

OpenStudy (anonymous):

(-1)^n.n^2=f(n)

OpenStudy (anonymous):

it ain't correct :D let me think

OpenStudy (anonymous):

so it he right

OpenStudy (anonymous):

-n^2=F(n) can be correct function

OpenStudy (anonymous):

what!!!!!!!!!!!!!!!! me dont get

OpenStudy (anonymous):

i think hes right @undeadknight26 @WinryRockbell @_hmm_

OpenStudy (anonymous):

u can check this F(n) function.

OpenStudy (anonymous):

what is F(n)

OpenStudy (anonymous):

function. n is input, -n^2 is output